OpenStudy (anonymous):

can you solve it The value of 1 -(2 -(3 -(.....100.....) is..........

OpenStudy (shubhamsrg):

100's the last term right ?

OpenStudy (anonymous):

hi shubham

OpenStudy (shubhamsrg):

is hi ka matlab main yes samjhun? :P

OpenStudy (anonymous):

okay

OpenStudy (phoenixfire):

Wouldn't you just work your way out from the inside. 3-100=-97 2-(-97)=99 1-99=-98

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

@PhoenixFire you are wrong! nice try

OpenStudy (phoenixfire):

How is it wrong?

OpenStudy (phoenixfire):

Oh.. wait

OpenStudy (phoenixfire):

Associative rule. it's the same as 1-2-3-100... work left to right. 1-2=-1 -1-3=-4 -4-100=-104

OpenStudy (anonymous):

for u r support I will give you options : A) - 40 B) - 50 c) 50 d) -100 e ) 100 f) 99

OpenStudy (anonymous):

It will be like: 1-2+3-4+5-6+7...-100

OpenStudy (anonymous):

Sum odd number from 1 to 100 - sum of even number from 1 to 100

OpenStudy (anonymous):

Still m not satisfied !! more proofs needed

OpenStudy (phoenixfire):

What do you mean by (......100.......)

OpenStudy (anonymous):

100

OpenStudy (anonymous):

yeah why there is (.....100...) ?

OpenStudy (phoenixfire):

so it's 1-(2-(3-(100))).... why did you add ........

OpenStudy (anonymous):

hmmmm..................

OpenStudy (anonymous):

you could write it like @PhoenixFire , what is the significance of those dots?

OpenStudy (anonymous):

for your knowledge its INEQUALITY QUESTIOn

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Sum of odd number from 1 to 100 -sum of even number from 1 to 100 2500-2550 -50

OpenStudy (anonymous):

A simple example : 1/2·3/4·5/6· ... ·99/100 < 1/10. solution : A(n) < 1/√3n,

hartnn (hartnn):

its just 1-2+3-4+5-6+......100 isn't it ?

hartnn (hartnn):

then @sauravshakya is right

OpenStudy (anonymous):

@sauravshakya you were correct but still add a thing and a medal for u is waiting! :p

OpenStudy (anonymous):

?

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

how can it be Sum of odd number from 1 to 100? there is -3 not 3

OpenStudy (anonymous):

1 -(2 -(3 -(.....100.....) =1-2+(3-(...100...) =1-2+3-(4-(5...100...)

OpenStudy (shubhamsrg):

lets start from inside,, ............-(95 - (96 - (97 - ( 98 - ( 99 -100) ) ) ) ) ) ............-(95 - (96 -(97 - (98 - 99 + 100 ) .... ..........-(95 - (96 - (97 - 98 + 99 - 100 ).. .........-(95 - 96 + 97 - 98 + 99 - 100) sign alternates everytime for each no..i.e. 2 consecutive nos. have opposite signs.. in the end, we shall have 1- 2 +3 - 4 ......... -100 (1+3+5......99) - (2 +4 +6 ...100) 25(100 - 102) =-50 was that corretc ?

OpenStudy (anonymous):

Open the bracket one by one and see

OpenStudy (anonymous):

YEp very correct @shubhamsrg

OpenStudy (anonymous):

good @shubhamsrg